OpenStudy (anonymous):

solve using the quadratic formula leave in simplified form 5x^2-3x-3=0 how would i do this?

OpenStudy (anonymous):

\[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \(a=5, b=-3, c=-3\)

OpenStudy (anonymous):

you get \[\frac{3\pm\sqrt{(-3)^2-4\times 5 \times (-3)}}{2\times 5}\] as a first step

OpenStudy (anonymous):

i got \[3\pm \sqrt{-51\div10}\]

OpenStudy (anonymous):

some mistake inside the radical

OpenStudy (anonymous):

\[\frac{3\pm\sqrt{9+60}}{10}\]

OpenStudy (anonymous):

oh right a - times a - is a positive \[3\pm \sqrt69/10\]

OpenStudy (anonymous):

that is it

OpenStudy (anonymous):

thats the quadratic formula? i know that what they give us os we have to plug it in to the original formula but the way u showed me is reallly simple thank you so so much :D

OpenStudy (anonymous):

yes, that is the quadratic formula don't forget that EVERYTHING is divided by \(2a\) not just the stuff in the radical

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

everything thats on top of it right

OpenStudy (anonymous):

yes, it is all on top of \(2a\) exactly as i wrote in the first post

OpenStudy (anonymous):

alright got it thanks a bunch once again :)

OpenStudy (anonymous):

yw (again)

OpenStudy (anonymous):

:D